ganeshie8 (ganeshie8):

Two identical cylindrical vessels with their bases at the same level each contain a liquid of density 1.30 × 10^3 kg/m^3. The area of each base is 4.00 cm^2, but in one vessel the liquid height is 0.854 m and in the other it is 1.560 m. Find the work done by the gravitational force in equalizing the levels when the two vessels are connected.

OpenStudy (amilapsn):

think about the initial and final positions of the center of gravity of the liquid.

ganeshie8 (ganeshie8):

I think the difference in y coordinate of center of gravity of the displaced liquid would be : (1.56-0.854)/2 = 0.353

ganeshie8 (ganeshie8):

perhaps it is 0.353/2 ?

OpenStudy (amilapsn):

think a bit more.... drawing a sketch would help

OpenStudy (amilapsn):

yep you're right. its .353/2

ganeshie8 (ganeshie8):

should i simply find the mass and calculate the reduced potential energy of that displaced liquid ?

OpenStudy (amilapsn):

yep!

ganeshie8 (ganeshie8):

4/10000*1.3*1000*9.8*0.353 = 1.799 J let me check the answer...

ganeshie8 (ganeshie8):

answer is 0.635 J let me double check my calculations..

OpenStudy (amilapsn):

work done=displaced liquid's mass\(\times\)g\(\times\)displacement of the center of gravity

ganeshie8 (ganeshie8):

mass = density * volume = 1.3 * 10^3 * (4/10000*0.353) = 0.184 kg

ganeshie8 (ganeshie8):

does that look good ?

OpenStudy (amilapsn):

yeah..

ganeshie8 (ganeshie8):

W = mgh = 0.184*9.8*0.353/2 = 0.318 J still no luck...

OpenStudy (amilapsn):

wait a minute.

ganeshie8 (ganeshie8):

okie

imqwerty (imqwerty):